# A Hessian for Gaussian Mixture Likelihoods in Nonlinear Least Squares

### v1

@@ Abstract
This paper proposes a novel Hessian approximation for Maximum a Posteriori estimation problems in robotics involving Gaussian mixture likelihoods. The proposed Hessian leads to better convergence properties. Previous approaches manipulate the Gaussian mixture likelihood into a form that allows the problem to be represented as a nonlinear least squares (NLS) problem. However, they result in an inaccurate Hessian approximation due to additional nonlinearities that are not accounted for in NLS solvers. The proposed Hessian approximation is derived by setting the Hessians of the Gaussian mixture component errors to zero, which is the same starting point as for the Gauss-Newton Hessian approximation for NLS, and using the chain rule to account for additional nonlinearities. The proposed Hessian approximation is more accurate, resulting in improved convergence properties that are demonstrated on simulated and real-world experiments. A method to maintain compatibility with existing solvers, such as ceres, is also presented. Accompanying software and supplementary material can be found at https://github.com/decargroup/hessian_sum_mixtures.
@@ I Introduction and Related Work
Estimating the state of a system from noisy and incomplete sensor data is a central task for autonomous systems. To describe inherent uncertainty in the measurements and state, probabilistic tools are required. Gaussian measurement likelihoods are commonly used in state estimation, allowing estimation problems to be easily formulated as instances of nonlinear least squares (NLS) optimization [1, §4.3]. The NLS formulations for state estimation stem from considering the negative log-likelihood (NLL) of the Gaussian, allowing a cancellation between the logarithm and the Gaussian exponent. However, in practice, many sensor models are highly non-Gaussian, as in underwater acoustic positioning [2], or subsea Simultaneous Localization and Mapping (SLAM) [3].
Even when the sensor characteristics are well-modeled by Gaussian distributions, multimodal error distributions arise in real-world situations involving ambiguity. Examples include simultaneous localization and mapping (SLAM) under unknown data associations [4], 3D multi-object tracking [5], and loop-closure ambiguity in pose-graph optimization [6]. In such cases, Gaussian measurement likelihoods are insufficient, and realizing robust and safe autonomy requires more accurate modeling of the underlying distributions.
@@
A popular parametric model used to represent multimodal distributions is the Gaussian mixture model (GMM), composed of a weighted sum of Gaussian components [7, §2.3.9]. Extensions of the incremental smoothing and mapping framework iSAM2, such as multi-hypothesis iSAM [8], have been developed to represent ambiguity in state estimation problems using GMMs. However, these approaches are incompatible with standard NLS optimization methods and require a dedicated solver. Hence, other research has focused on incorporating GMM measurement likelihoods for use in standard NLS optimization frameworks. In the GMM case, the NLL does not simplify since the logarithm is unable to cancel with the Gaussian exponents. One of the most widely used methods to overcome this is the Max-Mixture [9], which approximates the summation over Gaussians with a maximum operator. This extracts the dominant component of the mixture, reducing the problem back to a single Gaussian. This technique allows the problem to be cast into a NLS form at the cost of introducing additional local minima. The extension of Robust Incremental least-Squares Estimation (RISE) to non-Gaussian models [10] allows for error terms with arbitrary distributions to be cast into instances of NLS minimization. The application of the framework in [10] to GMMs error terms is presented in [11] and called the Sum-Mixture method. While the approach in [10] works well for its intended scope, its application to GMMs results in issues when iterative optimization methods are used.
@@
Common iterative methods for solving NLS problems, such as Gauss-Newton, rely on an approximation of the Hessian of the cost function. When the Sum-Mixture approach is used in conjunction with such optimization techniques, the resulting Hessian approximation is inaccurate, leading to degraded performance. This inaccuracy was pointed out in [11], but not explained theoretically. The paper [11] then proposes an alternative to the Sum-Mixture approach, termed the Max-Sum-Mixture approach, which is a hybrid between the Max-Mixture and Sum-Mixture approaches. The dominant component of the mixture is factored out, creating a Max-Mixture error term, while the remaining non-dominant factor is treated using the approach of [10]. The Hessian contribution for the non-dominant factor is inaccurate in a similar fashion as that of the Sum-Mixture, and improving on this aspect is the focus of this paper.
The contribution of this paper is a novel Hessian approximation for NLS problems involving GMM terms, which is derived by using the Gauss-Newton Hessian approximation for the component errors corresponding to each component of the GMM, and using the chain rule to take into account the additional nonlinearities. This approach is similar in spirit to how the robust loss nonlinearity is treated in deriving the Triggs correction proposed in [12] and used in the ceres library [13]. Traditional NLS solvers do not allow the user to specify a separate Hessian, only requiring an error definition and its corresponding Jacobian. Therefore, this paper also proposes a method of defining an error and Jacobian that uses the proposed Hessian while maintaining compatibility with the NLS solver.
The more accurate Hessian approximation is showcased in Figure 1 for a simple 1D example, and the improvement in the Hessian accuracy leads to improved convergence properties of gradient-based optimization methods. The benefits of the proposed approach are demonstrated on a toy problem, a simulated point-set registration problem [11][14] and on a SLAM problem with unknown data associations on the “Lost in the Woods” dataset [15]. The implementation is open-sourced at https://github.com/decargroup/hessian_sum_mixtures and uses a modified version of the navlie NLS solver [16].
@@ II-A Newton’s Method
Newton’s method is a local optimization method that iteratively minimizes a quadratic approximation to the loss function [1, §4.3.1]. Denoting the optimization variable as $\mbf{x}\in\mathbb{R}^{n}$ and the loss function to be minimized as $J(\mbf{x})$, Newton’s method begins with an initial guess $\mbf{x}^{(0)}$, and iteratively updates the optimization variable using
$\mbf{x}^{(i+1)}$ $=\mbf{x}^{(i)}+\alpha^{(i)}\Delta\mbf{x}^{(i)},$ (1)
where $\alpha^{(i)}$ is a user-defined step size and $\Delta\mbf{x}^{(i)}$ is the descent direction. Setting the gradient of the local quadratic approximation of the loss at $\mbf{x}^{(i)}$ to zero results in the descent direction being given by [17, §2.2]
@@
where $\frac{\partial J(\mbf{x})}{\partial\mbf{x}}$ and $\frac{\partial^{2}J(\mbf{x})}{\partial\mbf{x}\partial\mbf{x}^{\mathsf{T}}}$ are the Jacobian and Hessian respectively, and are evaluated at the current iterate $\mbf{x}^{(i)}$. The notation $\frac{\partial J(\mbf{x})}{\partial\mbf{x}^{\mathsf{T}}}=\left(\frac{\partial J(\mbf{x})}{\partial\mbf{x}}\right)^{\mathsf{T}}$ is used throughout this paper. The iterate $\mbf{x}^{(i)}$ is updated using (1) and the process is iterated to convergence. The main drawbacks of this method are the need to compute the Hessian, as well as the requirement for the Hessian to be positive definite. The Hessian is only guaranteed to be positive semidefinite for convex losses [18, §3.1.4], which are infrequent in robotics applications.
@@ II-B Gauss-Newton Method
The Gauss-Newton method is an optimization approach applicable to problems that have a NLS structure. The loss function is written as the sum of squared error terms,
@@
where $J_{i}(\mbf{x})$ denotes a single summand, and ${n_{f}}$ is the number of summands. The Jacobian and Hessian are computed using the chain rule,
@@ II-C Maximum a Posteriori Estimation
Maximum a Posteriori (MAP) estimation seeks to estimate the system state $\mbf{x}\in\mathbb{R}^{n}$ by maximizing the posterior probability [1, Sec. 3.1.2], which for Gaussian error models results in the optimization problem
@@ III Gaussian Mixture Error Terms
where $w_{k}$ is the weight of the $k$’th Gaussian component. Note that here the sum is over components, whereas in Sec. II the sums considered are over problem error terms, of which only a single one is now considered. To find an expression of the NLL for error models of the form (9), first let $\alpha_{k}$ be the weight normalized by covariance, written as
$\alpha_{k}$ $=w_{k}\det(\mbf{R}_{k})^{-1/2}.$ (10)
Then, dropping state- and covariance-independent normalization constants and using the change of variables (8) yields the expression for the corresponding summand given by
@@ III-A Max-Mixtures
The Max-Mixture approach, first introduced in [9], replaces the summation in (12) with a $\max$ operator, such that the NLL is given by
$-\log p_{\max}(\mbf{x})$ $=-\log\left(\max_{k}\alpha_{k}\exp\left(-\frac{1}{2}\mbf{e}_{k}(\mbf{x})^{\mathsf{T}}\mbf{e}_{k}(\mbf{x})\right)\right).$
@@ III-B Sum-Mixtures
A general method for treating non-Gaussian likelihoods in NLS is proposed in [10]. For a general non-Gaussian NLL $-\log p_{\textrm{non-gauss}}(\mbf{x})$, the corresponding NLS error $e_{\textrm{non-gauss}}(\mbf{x})$ is derived by writing the relationship
@@
where ${\gamma_{\text{{{SM}}}}}$ is a normalization constant chosen such that ${\gamma_{\text{{{SM}}}}}\geq\log p_{\textrm{non-gauss}}(\mbf{x})$ to ensure that the square root argument is always positive. In this case, the constant $c$ absorbs the constant offset that is added by the normalization constant. The expression (16) allows the error definition
@@
Ensuring positivity of the square root argument can be achieved by setting the normalization constant to ${\gamma_{\text{{{SM}}}}}=\sum_{k=1}^{n_{k}}\alpha_{k}$ [19].
@@ III-C Max-Sum-Mixtures
The Max-Sum-Mixture is introduced in [11], and is a hybrid between the Max-Mixture and Sum-Mixture approaches. The dominant component of the mixture is factored out such that the corresponding summand is given by
@@
The first term of (III-C) is analogous to the Max-Mixture case. The second term is treated using the Sum-Mixture approach,
$e_{2}(\mbf{x})$ $=\sqrt{2}\sqrt{-\log\frac{1}{{\gamma_{\text{{{MSM}}}}}}\sum_{k=1}^{n_{k}}\frac{\alpha_{k}}{\alpha_{{k}^{*}}}\exp(-\mbf{e}_{k}^{\mathsf{T}}\mbf{e}_{k}+\mbf{e}_{{{k}^{*}}}^{\mathsf{T}}\mbf{e}_{{k}^{*}})},$
where ${{\gamma_{\text{{{MSM}}}}}}$ is chosen as ${{\gamma_{\text{{{MSM}}}}}}={n_{k}}\max\frac{\alpha_{k}}{\alpha_{{k}^{*}}}+\delta$. The parameter $\delta>0$ is a damping constant [19] that dampens the influence of this nonlinear term in the optimization.
@@ IV Inaccuracy of the Non-Dominant Component Max-Sum-Mixture Hessian
NLS problems are defined by the choice of error. Different choices of error may be made, which in turn defines the Gauss-Newton Hessian approximation. Because of the Gaussian noise assumption in most robotics problems, the error choice is natural because the negative log-likelihood is quadratic. However, since GMMs do not yield a quadratic log-likelihood, there is no natural error choice. The Sum-Mixture is one such error choice but yields an inaccurate Hessian approximation. The inaccuracy is made clear by considering a problem with a single Gaussian mixture error term of the form (9), having a single mode. The cost function is given by
$J(\mbf{x})=\frac{1}{2}\mbf{e}(\mbf{x})^{\mathsf{T}}\mbf{e}(\mbf{x}),$ (21)
and when the Sum-Mixture formulation is applied to this problem, the cost function becomes
@@
with $e_{{\textrm{SM}}}(\mbf{x})=\sqrt{\mbf{e}^{\mathsf{T}}(\mbf{x})\mbf{e}(\mbf{x})}$. Even for a Gaussian likelihood problem that is easily solved by Gauss-Newton in its original formulation, since the error $e_{{\textrm{SM}}}(\mbf{x})$ is a scalar, the Sum-Mixture approach results in a rank one Gauss-Newton Hessian approximation for the corresponding summand. While this is not an issue within the original application in [10], the Hessian approximation when applied to mixtures is inaccurate.
@@
However, the non-dominant components in the second term of (III-C) are still subjected to the same treatment. Their contribution to the overall Hessian is thus inaccurate, leading to the discrepancies shown in Figure 1. In turn, the inaccurate Hessian negatively affects convergence, since it is used to compute the iteration step (2). Furthermore, it can affect consistency, since the Laplace approximation [7, Sec. 4.4] uses the Hessian to provide a covariance on the resulting state estimate.

### v2

# library.bib

@@ Abstract
This paper proposes a novel Hessian approximation for Maximum a Posteriori estimation problems in robotics involving Gaussian mixture likelihoods. Previous approaches manipulate the Gaussian mixture likelihood into a form that allows the problem to be represented as a nonlinear least squares (NLS) problem. The resulting Hessian approximation used within NLS solvers from these approaches neglects certain nonlinearities. The proposed Hessian approximation is derived by setting the Hessians of the Gaussian mixture component errors to zero, which is the same starting point as for the Gauss-Newton Hessian approximation for NLS, and using the chain rule to account for additional nonlinearities. The proposed Hessian approximation results in improved convergence speed and uncertainty characterization for simulated experiments, and similar performance to the state of the art on real-world experiments. A method to maintain compatibility with existing solvers, such as ceres, is also presented. Accompanying software and supplementary material can be found at https://github.com/decargroup/hessian_sum_mixtures.
Index Terms:
Sensor Fusion, Localization, Optimization and Optimal Control, Probabilistic Inference, SLAM
@@ I Introduction and Related Work
Estimating the state of a system from noisy and incomplete sensor data is a central task for autonomous systems. To describe inherent uncertainty in the measurements and state, probabilistic tools are required. Gaussian measurement likelihoods are commonly used in state estimation, allowing estimation problems to be easily formulated as instances of nonlinear least squares (NLS) optimization [barfootStateEstimationRobotics2017, §4.3]. The NLS formulations for state estimation stem from considering the negative log-likelihood (NLL) of the Gaussian, allowing a cancellation between the logarithm and the Gaussian exponent. Furthermore, the structure of NLS problems allows for the use of efficient solvers such as ceres [Agarwal_Ceres_Solver_2022], whereas more general optimizers or non-parametric methods [doucet2001sequential] may be more computationally demanding. Therefore, the NLS structure, which for many problems arises from Gaussianity of the error model, is highly desirable. However, in practice, many sensor models are highly non-Gaussian, as in underwater acoustic positioning [cheung2019non], or subsea Simultaneous Localization and Mapping (SLAM) [newman2003pure].
@@
Even when the sensor characteristics are well-modeled by Gaussian distributions, multimodal error distributions arise in real-world situations involving ambiguity. Examples include simultaneous localization and mapping (SLAM) with unknown data associations [doherty2020probabilistic], 3D multi-object tracking [poschmann2020factor], and loop-closure ambiguity in pose-graph optimization [lee2013robust]. In such cases, Gaussian measurement likelihoods are insufficient, and realizing robust and safe autonomy requires more accurate modeling of the underlying distributions. A popular parametric model used to represent multimodal distributions is the Gaussian mixture model (GMM), composed of a weighted sum of Gaussian components [bishopPatternRecognitionMachine2006, §2.3.9]. Extensions of the incremental smoothing and mapping framework iSAM2, such as multi-hypothesis iSAM [hsiao2019mh], have been developed to represent ambiguity in state estimation problems. However, these approaches are incompatible with standard NLS optimization methods and require a dedicated solver. Hence, other research has focused on incorporating GMM measurement likelihoods for use in standard NLS optimization frameworks. In the GMM case, the NLL does not simplify since the logarithm cannot cancel with the Gaussian exponents, meaning that GMMs are not directly compatible with NLS solvers. A common method to overcome this is the Max-Mixture, introduced by Olson et al. [olsonInferenceNetworksMixtures2013], which approximates the summation over Gaussians with a maximum operator, reducing the problem back to a single Gaussian. This technique allows the problem to be cast into an NLS form at the cost of introducing additional local minima. Rosen et al. [rosenRobustIncrementalOnline2013] introduce the extension of Robust Incremental least-Squares Estimation (RISE) to non-Gaussian models, which allows for error terms with arbitrary distributions to be cast into instances of NLS minimization. The application of the framework in [rosenRobustIncrementalOnline2013] to GMMs error terms is presented in Pfeifer et al. [pfeiferAdvancingMixtureModels2021] and called the Sum-Mixture method. While the approach in [rosenRobustIncrementalOnline2013] works well for its intended scope, its application to GMMs results in issues when iterative optimization methods are used.
@@
Common iterative methods for solving NLS problems, such as Gauss-Newton, rely on an approximation of the Hessian of the cost function. Accuracy of the Hessian approximation manifests itself in convergence of the algorithm, since it is used to compute the step in NLS algorithms. Furthermore, the Hessian approximation can affect consistency, since the Laplace approximation [bishopPatternRecognitionMachine2006, Sec. 4.4] uses the Hessian to provide a covariance on the estimate computed using the NLS algorithm.
When the Sum-Mixture approach is used in conjunction with such optimization techniques, the resulting Hessian approximation is inaccurate, leading to degraded performance. This inaccuracy is pointed out by Pfeifer et al. in [pfeiferAdvancingMixtureModels2021], but not explained theoretically. Pfeifer et al. [pfeiferAdvancingMixtureModels2021] then propose an alternative to the Sum-Mixture approach, termed the Max-Sum-Mixture approach, which is a hybrid between the Max-Mixture and Sum-Mixture approaches. The dominant component of the mixture is factored out, creating a Max-Mixture error term, while the remaining non-dominant factor is treated using the approach of Rosen et al. [rosenRobustIncrementalOnline2013]. The Hessian contribution for the non-dominant factor is inaccurate in a similar fashion as that of the Sum-Mixture, and improving on this aspect is the focus of this paper.
The contribution of this paper is a novel Hessian approximation for NLS problems involving GMM terms. The proposed Hessian approximation is derived by using the Gauss-Newton Hessian approximation for the component errors corresponding to each component of the GMM, and using the chain rule to take into account the additional nonlinearities. This approach is similar in spirit to how the robust loss nonlinearity is treated when deriving the Triggs correction proposed in [triggs2000bundle] and used in the ceres library [Agarwal_Ceres_Solver_2022]. The proposed Hessian approximation has better mathematical properties in the sense that all mixture components contribute in the same way to the Hessian, without treating the dominant mixture component in a different framework than the rest, unlike other state-of-the-art approaches. Improvements to Hessian accuracy can therefore be expected in cases where multiple mixture components overlap. Furthermore, traditional NLS solvers do not allow the user to specify a separate Hessian, only requiring an error definition and its corresponding Jacobian. Therefore, this paper also proposes a method of defining an error and Jacobian that uses the proposed Hessian while maintaining compatibility with the NLS solver.
The proposed Hessian approximation is showcased in Figure 1 for a 1D example, where it is visually more accurate compared to state of the art. Performance of the proposed method is validated on a toy problem, a simulated point-set registration problem [pfeiferAdvancingMixtureModels2021][Jian2011] and on a SLAM problem with unknown data associations on the “Lost in the Woods” dataset [lostInTheWoodsDataset]. The implementation is open-sourced at https://github.com/decargroup/hessian_sum_mixtures and uses the navlie NLS solver [cossette2023navlie].
@@ II-A Newton’s Method
Newton’s method is a local optimization method that iteratively minimizes a quadratic approximation to the loss function [barfootStateEstimationRobotics2017, §4.3.1]. Denoting the optimization variable as $\mbf{x}\in\mathbb{R}^{n}$ and the loss function to be minimized as $J(\mbf{x})$, Newton’s method begins with an initial guess $\mbf{x}^{(0)}$, and iteratively updates the optimization variable using $\mbf{x}^{(i+1)}=\mbf{x}^{(i)}+\alpha^{(i)}\Delta\mbf{x}^{(i)}$ where $\alpha^{(i)}$ is a user-defined step size and $\Delta\mbf{x}^{(i)}$ is the descent direction. Setting the gradient of the local quadratic approximation of the loss at $\mbf{x}^{(i)}$ to zero results in the descent direction being given by [madsen2004methods, §2.2]
@@
where $\frac{\partial J(\mbf{x})}{\partial\mbf{x}}{{\color[rgb]{0,0,0}\in\mathbb{R}^{1\times n}}}$ and $\frac{\partial^{2}J(\mbf{x})}{\partial\mbf{x}\partial\mbf{x}^{\mathsf{T}}}{{\color[rgb]{0,0,0}\in\mathbb{R}^{n\times n}}}$ are the loss Jacobian and Hessian respectively, and are evaluated at the current iterate $\mbf{x}^{(i)}$. The notation $\frac{\partial J(\mbf{x})}{\partial\mbf{x}^{\mathsf{T}}}=\left(\frac{\partial J(\mbf{x})}{\partial\mbf{x}}\right)^{\mathsf{T}}$ is used throughout this paper. The iterate $\mbf{x}^{(i)}$ is updated and the process is iterated to convergence. The main drawbacks of this method are the need to compute the Hessian, as well as the requirement for the Hessian to be positive definite. The Hessian is only guaranteed to be positive semidefinite for convex losses [boyd2004convex, §3.1.4], which are infrequent in robotics applications.
@@ II-B Gauss-Newton Method
The Gauss-Newton method is an optimization approach applicable to problems that have a NLS structure. This structure is leveraged to form an approximate Hessian that is in turn used in the Newton step (1). The loss function is written as the sum of squared error terms,
@@
where $J_{i}(\mbf{x})$ denotes a single summand, ${n_{f}}$ is the number of summands, and $\mbf{e}_{i}(\mbf{x})\in\mathbb{R}^{n_{e,i}}$. The Jacobian and Hessian are computed using the chain rule,
@@ II-C Maximum a Posteriori Estimation
Maximum a Posteriori (MAP) estimation seeks to estimate the system state $\mbf{x}\in\mathbb{R}^{n}$ by maximizing the posterior probability [barfootStateEstimationRobotics2017, Sec. 3.1.2], which for Gaussian error models results in the optimization problem
@@ III Gaussian Mixture Error Terms
where $w_{k}$ is the weight of the $k$’th Gaussian component. Note that here the sum is over components, whereas in Sec. II the sums considered are over problem error terms, of which only a single one is now considered. To find an expression of the NLL for error models of the form (8), first let $\alpha_{k}$ be the weight normalized by covariance, written as $\alpha_{k}=w_{k}\det(\mbf{R}_{k})^{-1/2}$. Then, dropping state- and covariance-independent normalization constants and using the change of variables (7) yields the expression for the corresponding summand given by
@@ III-A Max-Mixtures
The Max-Mixture approach, first introduced in Olson et al. [olsonInferenceNetworksMixtures2013], replaces the summation in (10) with a $\max$ operator, such that the NLL is given by
$-\log p_{\max}(\mbf{x})$ $=-\log\left(\max_{k}\alpha_{k}\exp\left(-\frac{1}{2}\mbf{e}_{k}(\mbf{x})^{\mathsf{T}}\mbf{e}_{k}(\mbf{x})\right)\right).$ (11)
@@ III-B Sum-Mixtures
A general method for treating non-Gaussian likelihoods in NLS is proposed in Rosen et al. [rosenRobustIncrementalOnline2013]. For a general non-Gaussian NLL $-\log p_{\textrm{non-gauss}}(\mbf{x})$, the corresponding NLS error $e_{\textrm{non-gauss}}(\mbf{x})$ is derived by writing the relationship
@@
where ${\gamma_{\text{{{SM}}}}}$ is a normalization constant chosen such that ${\gamma_{\text{{{SM}}}}}\geq p_{\textrm{non-gauss}}(\mbf{x})$ to ensure that the square root argument is always positive. In this case, the constant $c$ absorbs the constant offset that is added by the normalization constant. The expression (15) allows the error definition
@@
Ensuring positivity of the square root argument can be achieved by setting the normalization constant to ${\gamma_{\text{{{SM}}}}}=\sum_{k=1}^{n_{k}}\alpha_{k}$ [pfeifer2023adaptive].
@@ III-C Max-Sum-Mixtures
The Max-Sum-Mixture is introduced by Pfeifer et al. [pfeiferAdvancingMixtureModels2021], and is a hybrid between the Max-Mixture and Sum-Mixture approaches. The dominant component of the mixture is factored out such that the corresponding summand is given by
@@
The first term of (19) is analogous to the Max-Mixture case. The second term is treated using the Sum-Mixture approach,
$e_{2}(\mbf{x})$ $=\sqrt{2}\sqrt{-\log\frac{1}{{\gamma_{\text{{{MSM}}}}}}\sum_{k=1}^{n_{k}}\frac{\alpha_{k}}{\alpha_{{k}^{*}}}\exp(-\mbf{e}_{k}^{\mathsf{T}}\mbf{e}_{k}+\mbf{e}_{{{k}^{*}}}^{\mathsf{T}}\mbf{e}_{{k}^{*}})},$ (20)
where ${{\gamma_{\text{{{MSM}}}}}}$ is chosen as ${{\gamma_{\text{{{MSM}}}}}}={n_{k}}\max\frac{\alpha_{k}}{\alpha_{{k}^{*}}}+\delta$. The parameter $\delta>0$ is a damping constant [pfeifer2023adaptive] that dampens the influence of this nonlinear term in the optimization.
@@ IV Inaccuracy of the Non-Dominant Component Max-Sum-Mixture Hessian
NLS problems are defined by the choice of error. Different choices of error may be made, which in turn defines the Gauss-Newton Hessian approximation. Because of the Gaussian noise assumption in most robotics problems, the error choice is natural because the negative log-likelihood is quadratic. However, since GMMs do not yield a quadratic log-likelihood, there is no natural error choice. The Sum-Mixture is one such error choice but yields an inaccurate Hessian approximation. The inaccuracy is made clear by considering a problem with a single Gaussian mixture error term of the form (8), having a single mode. The cost function is given by $J(\mbf{x})=\frac{1}{2}\mbf{e}(\mbf{x})^{\mathsf{T}}\mbf{e}(\mbf{x})$. When the Sum-Mixture formulation is applied to this problem, the cost function becomes
@@
with $e_{{\textrm{SM}}}(\mbf{x})=\sqrt{\mbf{e}^{\mathsf{T}}(\mbf{x})\mbf{e}(\mbf{x})}$. Even for a Gaussian likelihood problem that is easily solved by Gauss-Newton in its original formulation, since the error $e_{{\textrm{SM}}}(\mbf{x})$ is a scalar, the Sum-Mixture approach results in a rank one Hessian approximation (5) for the corresponding summand. For state dimension $n>1$, the Hessian approximation is non-invertible and the Newton update (1) fails. While this is not an issue within the original application in [rosenRobustIncrementalOnline2013], the Hessian approximation when applied to mixtures is inaccurate as demonstrated with this base case.
@@
However, the non-dominant components in the second term of (19) are still subjected to the same treatment. Their contribution to the overall Hessian is thus inaccurate, leading to the discrepancies shown in Figure 1.